The average of three number is 20. If two number are 16 and 22.the third is?
Answer by Timnewman(323)

hi dear,
let the third number be x
since their average is 20;
Then,
(16+22+x)/3=20
(38+x)/3=20
38+x=60
x=60-38
x=22
The third number is 22